What is the mass of a 400- mL sample of ethyl alcohol if it’s density is 0.8 g/mL
Answer by math_tutor2020(3835)


density = mass/volume
d = m/v
d*v = m
m = d*v

Plug in the given volume and density
m = d*v
m = (0.8 g/mL)*(400 mL)
m = 320 grams

Note the mL unit cancels

Answer: 320 grams